What is the probability of obtaining tails six times in a row when flipping?
You have a 1 in 2 chance with one flip; so, 2 to the 6th power, or 1 in 2x2x2x2x2x2; a 1 in 64 chance.
How many different ways can you arranged the word china?
25 and over
-----------------------------------------------------------------------------------------------
The number of ways is 5x4x3x2x1= 120 ways
How are experimental probability and theorectical probability alike?
Both are measures of the likelihood of events.
What is the probability of the spinner landing on a vowel and then a consonant?
What is the probability of the spinner landing on CorB
The probability of drawing three black cards one at a time with replacement from a standard deck of 52 cards is 1/3x1/2x26/52, which is 0.833.
What percentage of the data falls outside 2 standard deviation of the mean?
4.55% falls outside the mean at 2 standard deviation
Which set of events is an example of dependent events?
A dependent event is one that relies on something else to happen.
For instance with a mouse trap the trap closing "depends" on the trigger being activated.
Use of probability in logarithmic relations?
None.
If you have an exact relationship - whether it is linear, polynomial, logarithmic or whatever - probability has no role to play.
How does a sampling error affect the interpretation of your data?
The greater the sampling error the greater the uncertainty about the results and therefore the more careful you need to be in the interpretation.
3 in 10
4
What strategy is used for multiple choice questions?
This one is for free: stay tuned at the end though.
These strategies only work if the test designer is not aware of the human behavior normal in test design. Being one of those people, my assessments were one third true/false divided 50/50; one third multiple choice of only four options and those were divided into 25% for A, B, C, & D; and one third essay.
Take notes and study.
What is the probability of getting green if there is 3 blue 2red and 4 green?
If there is 3 blue 2 red and 4 green. What is the probability of getting green?
It is 1/13*1/51 = 1/663 = 0.001508 approx.
25.00
The chance that a given event will occur usually expressed between the number 0 will not and 1 the event will occur
What is the probability of pulling a 5 or a heart from a deck of cards?
Well, if you include the jokers, you have 52 cards total. There are four suits, so that's four "5" cards in the deck, which gives you a 4 in 52 (or 2 in 21, simplified) probability. Each suit has 13 cards, so you have a 13 in 52 chance of pulling a card of any particular suit (13 is a prime number, and can't be simplified).